40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Sep 14 15:57:05 2021, jamie, Frogs, CDS, fb1 /var full after reboot, caused all sorts of problems 
    Reply  Tue Sep 14 16:44:54 2021, jamie, Frogs, CDS, fb1 /var full after reboot, caused all sorts of problems 
       Reply  Tue Sep 14 17:22:21 2021, Anchal, Update, CDS, Added temp sensor channels to DAQ list 
Message ID: 16330     Entry time: Tue Sep 14 17:22:21 2021     In reply to: 16327
Author: Anchal 
Type: Update 
Category: CDS 
Subject: Added temp sensor channels to DAQ list 

[Tega, Paco, Anchal]

We attempted to reboot fb1 daqd today to get the new temperature sensor channels recording. However, the FE models got stuck, apparantely due to reasons explaine din 40m/16325. Jamie cleared the /var/logs in fb1 so that FE can reboot. We were able to reboot the FE machines after this work successfully and get the models running too. During the day, the FE machines were shut down manually and brought back on manually, a couple of times on the c1iscex machine. Only change in fb1 is in the /opt/rtcds/caltech/c1/chans/daq/C0EDCU.ini where the new channels were added, and some hacking was done by Jamie in gpstime module (See 40m/16327).

ELOG V3.1.3-